Rahul's father is 3 times as old as Rahull. Four years ago, his father was 4 times older than him. How old is Rahul?
Answer by TimothyLamb(4379)

---
x = father
y = rahull
---
x = 3y
x - 4 = 4(y - 4)
---
put the system of linear equations into standard form
---
x - 3y = 0
x - 4 = 4y -16
---
x - 3y = 0
x - 4y = -12
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = father = 36
y = rahull = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php